4(x+1)^2 +2(y-2)^2 +(z-1)^2=1
1
Expert's answer
2020-12-23T18:51:41-0500


4(x+1)2 +2(y-2)2 +(z-1)2=1

Finding the equation of the tangent plane to the surface which passes through the point (-1,2,2)

Equation of the tangent plane is

Fx(xo,yo,zo)(x-xo) +Fy(xo,yo,zo)(y-yo) + Fz(xo,yo,zo)(z-zo)

xo=-1 ,yo=2 ,zo=2

F(x,y,z)=4(x+1)2 +2(y-2)2 +(z-1)2-1=0


Fx= 8x + 8 (i.e. partial derivative with respect to x)

Fx(-1,2,2)= 8(-1)+8=0


Fy= 4y - 8(i.e. partial derivative with respect to y)

Fy(-1,2,2)= 4(2) -8= 0


Fz= 2z - 2 (i.e. partial derivative with respect to z)

Fz(-1,2,2)= 2(2)-2=2


Equation of the tangent plane is

Fx(xo,yo,zo)(x-xo) +Fy(xo,yo,zo)(y-yo) + Fz(xo,yo,zo)(z-zo)=0

0(x+1) + 0(y-2) + 2(z-2)=0

2z - 4 =0

2z - 4 =0

The equation of the tangent that passes through the point (-1,2,2) is

2z - 4 =0